# Realistic Extreme Image Rescaling via Generative Latent Space Learning

### v1

@@ 6 Conclusion
Extreme image rescaling is a highly challenging yet significant task. In this paper, we propose a novel framework called LSBIR to address issues such as insufficient texture and semantic inaccuracies. Specifically, we leverage the rich priors encapsulated in the pre-trained SD model to mitigate the ill-posed nature of rescaling. In the first stage, we design a DFRM to achieve bidirectional mapping between HR features and LR images. In the second stage, the rescaled features from the first stage serve as conditions to guide the generation process of the SD model through an effective adapter. Both qualitative and quantitative experiments demonstrate the superiority of our approach.

### v2

@@ 7 Conclusion
We propose a novel framework called OSIRDiff to address issues such as insufficient texture and semantic inaccuracies in extreme image rescaling. Specifically, we design a DFRM to achieve bidirectional mapping between HR features and LR images. Then, we utilize a pre-trained SD model to perform a single denoising step on the rescaled features, enhancing its perceptual quality. Considering the non-uniform quality of the rescaled latent features, we propose a novel time-step alignment strategy to achieve a balance between fidelity and perceptual quality. Both qualitative and quantitative experiments demonstrate the superiority of our approach.